Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 70     Entry time: 30 Mar 2024, 22:50
Author: Chan Kim 
Group: SPB 
Subject: Shift summary 

X-ray delivery

- 12.72 keV, 10Hz, ~1.4 mJ/pulse, continuous beam drift in the horizontal direction observed

Optical laser delivery
- 487 nm ns laser for pump-probe
- 800 nm PP laser for jet illumination, stable delivery

Achievements/Observations
- Collected PP (pump-probe) dataset with 5 us time delay more than 12 hours
- Collected 3 runs (5 minutes run) of dark state in the end of the shift
- Jet was quite stable most of the time

Issues
- Continuous beam pointing drift in the horizontal direction observed
- Vacuum spike at MKB_PSLIT section trip the beam shortly which might be related to the beam pointing drift

ELOG V3.1.4-7c3fd00